OpenStudy (anonymous):

one number is 2 more than twice another. the sum of the number is 14. find the numbers?

OpenStudy (anonymous):

2 + 2x = 14 -2 -2 2x = 12 /2 /2 x = 6 I think that is the answer

OpenStudy (lilg132):

2x + 2 = 14